A particle moves with a uniform acceleration of 5m/s^2 in 10 seconds. what is the average distance covered in the last one second
1
Expert's answer
2019-06-03T10:11:59-0400

Ehe equatins of the movement with unifom acceleration:


"v=v_0+a*t"


"S=v_0*t+a*t^2\/2"

On the start of a movement:

"v_0=0, t=0"

Speed ("v_9" ) and distance ("S_9") after 9 seconds of movement:


"v_9=a*t_9; S_9=a*t_9^2\/2"

Where "t_9=9s" - time from movement start during 9 seconds

Distance ("S_{10}"​) after 10 seconds of movement:


"S_{10}=a*t_{10}^2\/2"

Where "t_{10}=10s" - time from bedinning during 10 seconds

The distance covered durind last decond:


"S_x=S_{10}-S_9=a*t_{10}^2\/2-a*t_9^2\/2"

Usind the numbers in SI:


"S_x=47.5m"

Answer: "S_x=47.5m"